OpenStudy (anonymous):

Evaluate the function f(p) 1/2p -2, when p=8

OpenStudy (anonymous):

Is it \(f(p)=\frac {1}{2p}-2\) or \(f(p)=\frac {1}{2p-2}\)

OpenStudy (anonymous):

No, f(p) 1/2 p -2, when p=8

OpenStudy (anonymous):

I mean 6 1/2